### v2

# A search for super-imposed oscillations to the primordial power spectrum in Planck and SPT-3G 2018 data

Abstract
We search for super-imposed oscillations linearly or logarithmically spaced in Fourier wavenumbers in Planck and South Pole Telescope (SPT-3G) 2018 temperature and polarization data. The SPT-3G temperature and polarization data provide a new window to test these oscillations at high multipoles beyond the Planck angular resolution and sensitivity. We consider both models with a constant and a Gaussian modulated amplitude, which correspond to three and four additional parameters beyond power-law primordial power spectrum for the templates considered, respectively. We find that each of the four models considered can provide an improved fit to Planck data, consistently with previous findings, and to SPT-3G data, always compared to power-law power spectrum. For a constant amplitude of the superimposed oscillations, we find tighter constraints on the amplitude of the oscillations from the combined Planck/SPT-3G data set than in each individual data sets. When the ranges of parameters which provide a better fit to Planck and SPT-3G data overlap, as in the case of Gaussian modulated oscillations, we find a larger $\Delta\chi^{2}\sim-17.5\,(-14.7)$ for logarithmic (linear) oscillations - in a combined Planck/SPT-3G data set than in each individual data sets. These findings will be further tested with upcoming CMB temperature and polarization measurements at high multipoles provided by ongoing ground experiments.
1 Introduction
[…]
Among the deviations from a power-law in the primordial spectrum, super-imposed oscillations, either linear or logarithmic in wavenumber, stand out due to their deeper theoretical foundations and improved fit to the Planck data. Non-vacuum initial vacuum states for quantum fluctuations [2, 3, 4], resonant models including periodic oscillations in the potential [5] as axion monodromy [6] are among the mechanisms which generate super-imposed oscillations. These models have therefore become a workhorse for exploration with CMB anisotropy data [1, 7, 8, 9, 10, 11, 12, 13, 14, 15], also by a joint power-spectrum/bispectrum analysis [16, 17], and with galaxy surveys data [18, 19, 20, 21]. These models constitute a benchmark for future CMB experiments [22, 23], as well as for future and ongoing galaxy surveys [24, 25, 26, 27, 28, 29, 30]. See [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44] for other searches of primordial features in CMB data.
CMB polarization anisotropies should be a refined test for primordial features, given the sharpness of the polarization transfer functions compared to those in temperature [45]. Whereas $Planck$ remains the main full sky data set until LiteBIRD [46], better data at higher resolution from ground experiments such as Atacama Cosmology Telescope (ACT) [47] and South Pole Telescope (SPT) [48] are already in the public domain. The high degree of consistency between $Planck$ and SPT-3G data has been discussed in [48]. We therefore explore oscillations linear and logarithmically spaced in $k$ with a constant and a Gaussian modulated amplitude with $Planck$ [49] and, for the first time, with SPT-3G 2018 data [48], as well with their combination  [^1] See [43] for a study of different primordial features with $Planck$ and ACT DR4 data [47]. Note however that $Planck$ and ACT DR4 data lead to discrepant estimates for $\Omega_{b}h^{2}$ and $n_{\mathrm{s}}$ [47]..
This paper is structured as follows. After the introduction, in Section 2 we describe the four models chosen as representative for super-imposed oscillations, the data sets used and the sampling methodology. In Section 3, we describe our results and we conclude in Section 4.
2 Methodology
In this paper, we investigate the effect of sinusoidal oscillations in the primordial power spectrum which are periodic on linear or logarithmic scales. In our study, we further examine two class of models: one where the oscillations span the entire observable range of wavenumbers ($k$), and another where the oscillations are localized. In the latter case, these oscillations undergo damping on both sides of a central peak. In total, we analyse four models in this study: linear and logarithmic oscillations, with constant amplitude and Gaussian modulated amplitude, as can be seen in Fig. 1
Figure 1: Primordial curvature power spectra for representative models of the four templates considered.
2.1 Models and Priors
We consider the following templates for global oscillations:
$P_{\rm global}^{\rm lin}(k)=P_{0}(k)\left[1+\alpha\cos\left(\omega\,\frac{k}{k_{*}}+\phi\right)\right]$ (2.1)
$P_{\rm global}^{\rm log}(k)=P_{0}(k)\left[1+\alpha\cos\left(\omega\ln\frac{k}{k_{*}}+\phi\right)\right]$ (2.2)
These oscillations have an amplitude $\alpha$, frequency $\omega$ and a phase $\phi$ with priors listed in Table 1. The priors are the same of those employed in Ref. [1]. Here $P_{0}(k)$ is the power-law power spectrum given as $A_{s}(k/k_{*})^{n_{s}-1}$ where $k_{*}=0.05\,{\rm Mpc}^{-1}$.
As localized oscillation models extending Eqs.  2.12.2, we consider :
$P_{\rm local}^{\rm lin}(k)=P_{0}(k)\bigg{[}1+\alpha\cos\left(\omega\,\frac{k}{k_{*}}+\phi\right)e^{-\frac{\beta^{2}(k-\mu)^{2}}{2k_{*}^{2}}}\bigg{]}$ (2.3)
$P_{\rm local}^{\rm log}(k)=P_{0}(k)\bigg{[}1+\alpha\cos\left(\omega\ln\frac{k}{k_{*}}+\phi\right)e^{-\frac{\beta^{2}(k-\mu)^{2}}{2k_{*}^{2}}}\bigg{]}$ (2.4)
that contain five parameters, two, i.e. the center $\mu$ and the width $\beta$ of the Gaussian envelope, in addition to our global oscillation model. In this work we consider varying all the five parameters with priors as in Table 1. We have explicitly checked that the dimensional reduction adopted in [43], by fixing the phase in such a way that the peak of the Gaussian is located at a maximum of the oscillations $\phi=-\omega\mu/k_{*}$ ($\phi=-\omega\ln(\mu/k_{*})$) for the linear (logarithmic) case, modifies the results in the following way. The distribution of $\chi^{2}$ and the marginalized posterior for the amplitude are quite robust to fixing the phase. The Bayesian ratio, instead, would be different because of the smaller prior volume when fixing the phase. The prior ranges for these five parameters are listed in Table 1. Note that the priors chosen for $\alpha\,,\omega\,,\phi$ are the same as those of of the global oscillations case.
$\alpha$ $\log_{10}\omega$ $\mu$ $\beta$ $\phi/(2\pi)$
Linear [0, 0.5] [0,2.1] - - [0,1]
Log [0, 0.5] [0,2.1] - - [0,1]
Damped Linear [0, 0.5] [0,2.1] [0.001, 0.175] [0,10] [0,1]
Damped Log [0, 0.5] [0,2.1] [0.001, 0.175] [0,10] [0,1]
Table 1: Prior table for parameter of the super-imposed oscillation templates.
The priors for the six remaining cosmological parameters are the standard priors used in Polychord [50] and are given in Table 2. These priors are used for each of the four templates.
$\Omega_{\mathrm{b}}h^{2}$ $\Omega_{\mathrm{c}}h^{2}$ $100\theta_{\mathrm{M}C}$ $\tau$ $n_{\mathrm{s}}$ $\log[10^{10}A_{\mathrm{s}}]$
Prior [0.019, 0.025] [0.095,0.145] [1.03,1.05] [0.01,0.4] [0.885,1.04] [2.5,3.7]
Table 2: Prior table for cosmological parameters. Note that the priors on the background parameters are narrower than the standard MCMC analyses performed with $Planck$ data. For nested sampling we use this restricted prior range throughout our analyses as suggested in the PolyChord sampler [51]
2.2 Datasets and Samplers
Our analyses in this paper are based only on CMB data. We work with 3 combinations of datasets, namely $Planck$ (P18TP), SPT-3G (SPT3G) and combined P18TP and SPT-3G (P18TP+SPT3G).
We use $Planck$ 2018 release temperature and polarization likelihoods [49]. Following the primordial feature analysis of $Planck$ [1], we do not consider $Planck$ lensing likelihood. Differently from the $Planck$ analysis for features [1], we restrict ourselves to binned data, but all foreground and calibration nuisance parameters are allowed to vary along with their recommended priors [49]. In this paper we consider Plik binned data since the SPT-3G data are provided in a binned format, inhibiting the joint exploration of the high frequency oscillations domain allowed by unbinned Planck data.
For SPT-3G 2018 data, we use temperature, polarization auto and cross correlation data from 90, 150 and 220 GHz channels  [^2] https://pole.uchicago.edu/public/data/balkenhol22/ [48]. When we use SPT only likelihood for analysis, in order to break the degeneracy between the optical depth and the amplitude of primordial fluctuations, we use the optical depth prior from $Planck$ 2018 baseline analysis, $\tau=0.054\pm 0.074$, as recommended in [48]. Here, too, we allow the default nuisance parameters to vary.
The consistency and complementarity of $Planck$ 2018 and SPT-3G 2018 data is shown in Fig. 2.
Figure 2: $Planck$ (black) and SPT-3G TT, TE, EE (orange) binned residuals to the $Planck$ 2018 TP power law best-fit. The relative difference of the power law bestfit to SPT-3G data plus $Planck$ 2018 measurement of $\tau$ with respect to the $Planck$ 2018 TP power law bestfit is also shown for comparison (blue line).
Given the lack of publicly available covariance between $Planck$ and SPT-3G data, we consider as baseline a conservative combined $Planck$/SPT-3G data set in which there is no-overlap in multipoles between the two datasets. In gluing the two data sets, we follow indications in the SPT-3G release [48], where it has been discussed that SPT-3G data have smaller uncertainties than $Planck$ above multipole $\ell=2000$ for TT, $\ell=1400$ for TE and $\ell=1000$ for EE. Following that, we use a truncated version of $Planck$ Plik TTTEEE likelihood with TT in the range $\ell=30-2000$, TE in $\ell=30-1400$ and EE in $\ell=30-1000$. Therefore, in this combined data set, we use TT from $\ell=2000$, TE from $\ell=1400$ and EE from $\ell=1000$ for SPT-3G data. Note that with this conservative choice, the multipole range $10^{3}\lesssim\ell\lesssim 2\times 10^{3}$ are tested by $Planck$ in TT and by SPT-3G in EE.
[…]
Given the multi-modal nature of posteriors, Markov Chain Monte Carlo (MCMC) is not expected to be effective in sampling the parameter space. We therefore use the Polychord [50] sampler for nested sampling of parameter spaces and for the computation of Bayes factors. We use $n_{\rm live}=2048$ for all analyses ($n_{\rm live}=2000$ was also in used in Ref. [1] for the search of features). In settling for live points related to combined analysis, we compared MCMC results with Polychord for a power-law spectrum. For $n_{\rm live}=2048$ parameter posteriors obtained from Polychord converge to MCMC posteriors. We noticed that for the feature models too, $n_{\rm live}=2048$ is a conservative choice for the convergence of the posteriors. In obtaining best-fits, we use BOBYQA minimizer [53]. We identify the samples from the Polychord chains corresponding to maximum likelihood peak and we obtain local minima by BOBYQA around that peak.
Models $\Delta\chi^{2}_{\mathrm{P18}}$ $\ln B_{\mathrm{P18}}$ $\Delta\chi^{2}_{\mathrm{SPT}}$ $\ln B_{\mathrm{SPT}}$ $\Delta\chi^{2}_{\mathrm{P18+SPT}}$ $\ln B_{\mathrm{P18+SPT}}$
Lin undamped -11.8 -2.6 $\pm$ 0.3 -7.0 -1.8 $\pm$ 0.3 -12.0 -4.5 $\pm$ 0.4
Log undamped -9.3 -2.9 $\pm$ 0.3 -12.0 -1.2 $\pm$ 0.3 -14.3 -6.0 $\pm$ 0.4
Lin damped -11.8 -2.0 $\pm$ 0.3 -7.7 0.0 $\pm$ 0.3 -14.7 -1.8 $\pm$ 0.3
Log damped -10.0 -1.8 $\pm$ 0.3 -12.0 -0.4 $\pm$ 0.3 -17.5 -3.1 $\pm$ 0.4
Table 3: The table illustrates improvement in fit to the data and Bayesian evidence obtained by the templates with respect to the power law primordial power spectrum for the three data sets considered. We have studied the templates, both independently and jointly, against $Planck$ and SPT dataset. In all cases we get a good improvement in fit to the data with respect to power law power spectrum, although none of the models are preferred in terms of Bayesian evidence because of the number of additional parameters.
Figure 3: Marginalized joint 68 %, 95 % and 99 % regions for $(\alpha,\omega)$ for linear (left) and logarithmic undamped wiggles. Our results are quite consistent with those in Ref. [1], which are instead obtained with unbinned data and fixed nuisance/foreground parameters.
Figure 4: Undamped power spectrum templates: marginalized 68 % and 95 % posterior distribution for amplitude, frequency and phase of linear (on the left) and logarithmic templates (on the right) in Eqs 2.12.2. We study the templates using $Planck$ (P18TP), SPT (SPT3G) data individually and also with a non overlapping combination of both dataset (P18TP + SPT3G).
Figure 5: Damped power spectrum templates: marginalised 68 % and 95 % posterior distributions for amplitude, frequency and phase of linear (on the left) and logarithmic templates (on the right) present in Eqs 2.32.4. We study the templates using Planck (P18TP), SPT (SPT3G) data individually and also with a non overlapping combination of both dataset (P18TP + SPT3G). .
3 Results
Results for $Planck$, SPT-3G and the baseline combination of $Planck$ and SPT-3G data are presented in Table 3.
We first discuss the results obtained for the undamped templates as given in Eqs. 2.12.2. We find a very good agreement with the results in [1] when using $Planck$ 2018 unbinned data and nuisance/foreground parameters fixed to the power lawbestfits. The joint regions we obtain in Fig. 3 are quite similar in shape to those Ref. [1], but broader because we use a binned likelihood and the nuisance/foreground parameters are allowed to vary. Linear and logarithmic undamped templates lead to an improvement $\Delta\chi^{2}_{\mathrm{P18}}\sim-11.8$ and $\sim-9.3$, respectively, compatible with previous results [1, 54].
With SPT-3G 2018 and the $Planck$ prior on $\tau$, we find a prominent peak in the posterior for $\log_{10}\omega\sim 1.5$ with $\Delta\chi^{2}_{\mathrm{SPT}}\sim-12.0$ for logarithmic oscillations, whereas we find a smaller improvement in the fit for linear oscillations, always compared to power-law power spectrum. For SPT3G data, we find that a contribution to this improvement comes from fitting logarithmic oscillations in the range $1000\lesssim\ell\lesssim 2000$.
For these global oscillation models, the posteriors peak at different frequency for $Planck$ and SPT-3G 2018 data. For the $Planck$ + SPT3G data set, we find a $\chi^{2}_{\mathrm{P18+SPT}}\sim-12\,(-14.3)$ for linear (logarithmic) undamped oscillations. As is clear from Fig. 4, this combined $Planck$ + SPT-3G data set constrains the amplitude of these oscillations in a tighter way compared to the individual data sets. We obtain $\alpha\lesssim 0.031$ ($\alpha\lesssim 0.023$) at 95 % CL from $Planck$ plus SPT-3G for linear (logarithmic) oscillations, which improve on the corresponding $Planck$ 95 % CL constraint $\alpha\lesssim 0.036$ ($\alpha\lesssim 0.031$). None of the two models are statistically preferred over the power law model for any data set considered.
We analyzed the damped templates in Eqs. 2.32.4. SPT-3G data can constrain the envelope of these templates at scales beyond the $Planck$ resolution as the one mimicking the $A_{\rm L}$ effect. The triangle plots for the damped linear and logarithmic template corresponding to all three datasets are provided in Fig. 5. We find that the Gaussian modulated linear oscillations provide a larger improvement in the fit to $Planck$ data ($\Delta\chi^{2}_{\mathrm{P18}}\sim-11.8$) than Gaussian modulated logarithmic oscillations ($\Delta\chi^{2}_{\mathrm{P18}}\sim-10$). For SPT-3G data, the opposite occurs, i.e. Gaussian modulated logarithmic oscillations provide a better fit compared to linear oscillations (-12.0 vs -7.7). For a Gaussian modulated amplitude, we find a prominent peak at $\ln\omega_{\rm log}\sim 1.55$ for the SPT-3G data set (combined with the prior on $\tau$ from Planck 2018). These oscillations are related, but do not correspond exactly, to oscillations in SPT-3G EE data found by Gaussian regression in [55] [^3] Oscillations in $C_{\ell}^{EE}$ are not necessarily fit by super-imposed oscillations to the curvature power spectrum which are also imprinted in $C_{\ell}^{TT}$.. Since the ranges of parameters which provide a better fit independently to $Planck$ and SPT-3G data approximately overlap for modulated oscillations, we find a larger $\Delta\chi^{2}$ in the combined $Planck$ plus SPT-3G data set, up to the maximum $\Delta\chi^{2}\sim-17.5(-14.7)$ for modulated logarithmic (linear) oscillations. It is interesting to note that SPT-3G data do not prefer $A_{L}$ larger than 1 [48], but does not disfavour the modulated linear oscillation template which mimics the $A_{L}$ effect [1].
We would like to note that while the P18+SPT-3G 2D marginalized contours are tighter than both P18 and SPT-3G constraints, the constraints on the amplitude get worse compared to P18. This degradation can be explained by the positive correlation $\alpha-\mu$ in both plots. A higher value of amplitude of super-imposed oscillations is allowed when the feature is located at very small scales $\mu\gtrsim 0.1~{}\rm{Mpc}^{-1}$. In the joint likelihood, scales smaller than this wavenumber are not well constrained by $Planck$. Therefore, majority of the samples at $\mu\gtrsim 0.1~{}\rm{Mpc}^{-1}$ fit the SPT3G data better with the super-imposed oscillations while within the $Planck$ observed scales, the spectrum stays power law. For undamped cases, oscillations are present at all cosmological scales, we find that the joint dataset constrains the amplitude better than any individual datasets.
Figure 6: Relative differences of the $Planck$ 2018 TP best fits for the four models considered with respect to the corresponding $Planck$ 2018 TP power law bestfit (see legends for the colors). Plik TT, TE, EE binned residuals (black data points) to the $Planck$ 2018 TP power law best-fit are also plotted for comparison.
Figure 7: Top panel: relative differences of bestfits to the SPT-3G plus $Planck$ 2018 measurement of $\tau$ for the four models considered with respect to the corresponding SPT-3G power law bestfit (see legends for the colors). SPT-3G TT, TE, EE binned residuals (orange data points) to the SPT-3G power law best-fit are also plotted for comparison. Bottom panel: the same for $Planck$ combined with SPT-3G data.
In Fig. 6 we compare the relative differences of the $Planck$ 2018 TP bestfits for the four models considered with respect to the corresponding $Planck$ 2018 TP power law bestfit with the Plik TT, TE, EE binned residuals. In Fig. 7 we repeat this plot for SPT-3G and the $Planck$ plus SPT-3G combination.
4 Conclusions
We have searched for super-imposed oscillations linearly or logarithmically spaced in Fourier wavenumbers in Planck and SPT-3G 2018 temperature and polarization data. We study these super-imposed oscillations either with a constant amplitudes or amplitudes modulated with a Gaussian envelope; to our knowledge it is the first time that the modulated logarithmic oscillation is tested with current CMB data. SPT-3G, with better angular resolution and sensitivity can test primordial features at multipoles beyond the $Planck$ resolution. This is particularly advantageous for features that are at the edge of the $Planck$ angular resolution as the one mimicking the $A_{\rm L}$ effect. Cosmological parameters estimated from SPT-3G 2018 temperature and polarization data [48] in power law $\Lambda$CDM model are completelely consistent with $Planck$, and therefore, can be used in combination with $Planck$ to provide a precision test for models beyond the power law model such as those including features in the primordial power spectrum. See [48, 56, 57] for other studies of models beyond power law with SPT-3G data.
In this analysis, we use SPT-3G public (binned) data. We therefore use also $Planck$ binned 2018 data, differently from the search for features performed in Ref. [1], which used unbinned data but fixed nuisance and foreground parameters to their baseline values. We use PolyChord [50] as a sampler given the multimodal nature of the resulting posterior distributions and we allow for variation of the foreground and nuisance parameters.
For $Planck$, we find results consistent with those obtained in [1] for oscillations with a constant amplitude. When these oscillations have a Gaussian modulation, linear and logarithmic oscillations produce $\sim-11.8$ and $\sim-10.0$ as improvement in $Planck$ 2018 $\Delta\chi^{2}$, respectively.
We have tested these models with primordial oscillations with SPT-3G 2018 data for the first time. For SPT-3G 2018 data, we have found that all the models studied here improve the fit compared to the power law model, although they are not statistically preferred over the power-law power spectrum. As expected, for $\Lambda$CDM+power law cosmology [48], SPT-3G is less constraining than $Planck$ 2018 data. Our analysis shows that the frequencies preferred by SPT-3G 2018 data do not coincide with those preferred by $Planck$ 2018 data when the amplitude of the super-imposed oscillations is constant. However, for a Gaussian modulated amplitude, we find a prominent peak at $\log_{10}\omega_{\rm log}\sim 1.6$ and a broader peak at $\log_{10}\omega_{\rm lin}\gtrsim 1.1$. In particular, we find that super-imposed logarithmic oscillations lead to $\Delta\chi^{2}\sim-12$. These oscillations do not correspond exactly, but are related to certain features in SPT-3G EE data found by Gaussian regression in [55]  [^4] Oscillations in $C_{\ell}^{EE}$ are not necessarily fit by super-imposed oscillations to the curvature power spectrum which are also imprinted in $C_{\ell}^{TT}$..
We have therefore constrained these models with the combination of $Planck$ and SPT-3G data for the first time. Since a covariance between $Planck$ and SPT-3G data is not available, we consider the main combined data set in which we cut $Planck$ (SPT) above (below) the multipoles at which the $Planck$ S/N is reached by SPT as stated in [48] and described in section 2. This conservative procedure which enforces just one measurement for each multipole bin should overcome the issue of potentially double counting information with a full $Planck$ and SPT-3G combination without a covariance matrix between the two experiments. By adopting this combination, we find that SPT-3G $-\ell$ polarization data play an important role for all the models studied.
We find that superimposed oscillations with a constant amplitude do not improve the fit and are tighter constrained by $Planck$+SPT-3G 2018 compared to $Planck$ 2018 data. When the ranges of parameters which provide a better fit independently to $Planck$ and SPT-3G data overlap, we find a larger $\Delta\chi^{2}$ in the combined $Planck$ SPT-3G data set, $\Delta\chi^{2}\sim-17.5$ (-14.7) for modulated logarithmic (linear) oscillations. It is interesting to note that SPT-3G data do not prefer $A_{L}$ larger than 1 [48], but do not disfavour the modulated linear oscillation template which mimics the $A_{L}$ effect [1]. This different response of SPT-3G data to different extensions of the power law model might be due to the number of extra parameters, 1 for $A_{L}$ and 5 for the modulated oscillation template.
Our findings could be extended to higher frequencies with data which have a finer binning and will be further tested by upcoming CMB temperature and polarization measurements at high multipoles by ongoing and future ground experiments, such as ACT, SPT, Simons Observatory [58] and CMB-S4 [59].
Acknowledgments
The authors acknowledge the use of computational resources at the Institute of Mathematical Science’s High Performance Computing facility [Nandadevi] and of CNAF HPC cluster in Bologna. We would like to thank Jan Hamann for useful comments on the manuscript. FF and DKH would like to thank Karim Benabed and Silvia Galli for discussions on SPT-3G 2018 likelihood. AA is supported by an appointment to the JRG Program at the Asia Pacific Center for Theoretical Physics through the Science and Technology Promotion Fund and Lottery Fund of the Korean Government, and was also supported by the Korean Local Governments in Gyeongsangbuk-do Province and Pohang City. AA also acknowledges support from the NRF of Korea (Grant No. NRF-2022R1F1A1061590) funded by the Korean Government (MSIT). AA, FF, DKH acknowledge travel support through the India-Italy "RELIC - Reconstructing Early and Late events In Cosmology" mobility program. FF would like to thank IMSc for warm hospitality for the December 2022 visit when this work started. FF, DP acknowledge financial support from the contract by the agreement n. 2020-9-HH.0 ASI-UniRM2 “Partecipazione italiana alla fase A della missione LiteBIRD". FF, DKH, DP acknowledge financial support from Progetti di Astrofisica Fondamentale INAF 2023. DKH would like to acknowledge the support from the Indo-French Centre for the Promotion of Advanced Research – CEFIPRA grant no. 6704-4. AS would like to acknowledge the support by National Research Foundation of Korea NRF-2021M3F7A1082056, and the support of the Korea Institute for Advanced Study (KIAS) grant funded by the government of Korea.